# 11footnotetext: Implementation of AQUA is public available at https://github.com/tychenn/AQUA

### v1

@@ Appendix C More Results of Harmlessness of AQUA
LLaVA-NeXT Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
@@
InternVL3 Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
@@
Qwen-VL-Chat Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
@@
Qwen2.5-VL-Instruct Acronym-replace $5.00$ $100\%$ $5.00$ $100\%$
@@
Appendix D More Results of Robustness of AQUA
In this section, we will test the robustness of our AQUA on more models in Table 8. Because robustness has little to do with the dataset, here we show the experimental results on the MMQA dataset as in the main text.
Table 8: This table shows the Rank and p-value of watermark image after some common transformations.
Models Transformations $\textit{AQUA}_{acronym}$ $\textit{AQUA}_{spatial}$
Rank $\downarrow$ p-value $\downarrow$ Rank $\downarrow$ p-value $\downarrow$
LLaVA-NeXT Rescale $1.03$ $7.57e^{-274}$ $1.36$ $3.62e^{-62}$
Rotate $1.07$ $3.76e^{-142}$ $1.61$ $3.41e^{-45}$
Gaussian $1.07$ $4.10e^{-264}$ $1.46$ $1.64e^{-56}$
Rescale + Rotate + Gaussian $1.06$ $1.70e^{-256}$ $1.79$ $1.43e^{-39}$
InternVL3 Rescale $1.03$ $4.85e^{-231}$ $1.36$ $4.85e^{-46}$
Rotate $1.07$ $9.17e^{-112}$ $1.61$ $1.59e^{-31}$
Gaussian $1.07$ $7.69e^{-229}$ $1.46$ $3.76e^{-39}$
Rescale + Rotate + Gaussian $1.06$ $6.04e^{-216}$ $1.79$ $8.24e^{-27}$
Qwen-VL-Chat Rescale $1.03$ $4.58e^{-154}$ $1.36$ $4.15e^{-45}$
Rotate $1.07$ $8.08e^{-106}$ $1.61$ $6.82e^{-32}$
Gaussian $1.07$ $5.26e^{-142}$ $1.46$ $9.14e^{-38}$
Rescale + Rotate + Gaussian $1.06$ $5.12e^{-125}$ $1.79$ $6.78e^{-25}$
Qwen2.5-VL-Instruct Rescale $1.03$ $3.19e^{-269}$ $1.36$ $7.68e^{-68}$
Rotate $1.07$ $7.54e^{-238}$ $1.61$ $4.15e^{-51}$
Gaussian $1.07$ $1.23e^{-251}$ $1.46$ $8.48e^{-62}$
Rescale + Rotate + Gaussian $1.06$ $5.47e^{-268}$ $1.79$ $6.49e^{-46}$

### v2

@@ C.2 More Results of Harmlessness of AQUA
LLaVA-NeXT Acronym-replace $10.00$ $\% 10.0010$.00 $100\%$
@@
InternVL3 Acronym-replace $10.00$ $\% 10.0010$.00 $100\%$
@@
Qwen-VL-Chat Acronym-replace $10.00$ $\% 10.0010$.00 $100\%$
@@
Qwen2.5-VL-Instruct Acronym-replace $10.00$ $\% 10.0010$.00 $100\%$
@@
C.3 More Results of Stealthiness of AQUA
Figure 7 illustrates the retrieval probability of watermarked images as a function of the number of injected images.
More Stringent VLM-based Filters Furthermore, we explored the use of other prompts to filter watermarked images from the AQUA method.
To evaluate a more stringent adaptive attack scenario for the AQUA method, we simulated a stricter OCR-based filter. In contrast to the primary approach where a VLM assesses 5 images collectively, this filter evaluates images individually. An image is flagged if the VLM detects any presence of text within it. Intuitively, this method is expected to screen out the majority of $\textbf{AQUA}_{acronym}$’s text-based watermarks, but at the cost of an exceptionally high FPR on benign images. The prompt used and the experimental results are presented below:
OCR Filter Simulator You are an image filter. You need to identify whether there is text in this picture. If yes, answer ’yes’, if not, answer ’no’.
Similarly, we conducted an experiment to assess the VLM’s ability to discern plausibility from its internal knowledge. Single image is input to the model, which is prompted to determine whether the content is unnatural or semantically incongruous. The prompt and corresponding results are detailed below:
Strange Image Filter from VLM’s Perspective You are an image filter. If you think this picture is unreasonable or unnatural, answer ’yes’, otherwise answer ’no’.
Table 12: FPR represents the proportion of non-watermarked images incorrectly identified as watermarked, while TPR is the proportion of watermarked images that are correctly identified.
Methods Metrics Exp. 1 Exp. 2 Exp. 3
$\textbf{AQUA}_{acronym}$ FPR $67.60\%$ $61.44\%$ $72.02\%$
TPR $\% 100100\%$ $100\%$
$\textbf{AQUA}_{spatial}$ FPR $2.76\%$ $3.44\%$ $4.10\%$
TPR $2\%$ $4\%$ $2\%$
Our experimental results indicate that while a stringent VLM-based filter can remove a subset of the watermarked images, it concurrently incurs a prohibitively FPR on benign images. Given that an adaptive adversary’s primary objective is to augment their RAG system’s capabilities with the unauthorized database, adopting a filter that severely degrades the quality of legitimate data is an impractical strategy. This operational constraint for the adversary further underscores the stealthiness of our AQUA watermarking framework.
C.4 More Results of Robustness of AQUA
In this section, we provide a comprehensive evaluation of AQUA’s robustness. We examine two critical aspects: (1) resilience against adaptive adversaries who attempt to circumvent the watermark through model refusals or adversarial fine-tuning, and (2) stability under benign system updates, including component replacements and dataset drifts.
C.4.1 Defense against Adversarial Adaptations
We first address the concern of model-level adaptations by an adversary aiming to erase or evade the watermark triggers.
Evaluation of Potential Refusals.
To empirically address the concern that advanced VLMs might perceive watermarked images as “unnatural” and consequently refuse to generate responses, we conducted direct evaluations using state-of-the-art models, including Gemini-3-Pro, GPT-5.1-High, and Qwen3-VL. In our experiments, we observed no refusal behaviors. These models consistently processed the spatial queries as standard visual reasoning tasks, confirming that the watermark triggers remain compatible with high-level visual perception and do not trigger safety refusals.
Robustness against Adversarial Fine-tuning.
We simulated an adaptive adversary who attempts to bypass the watermark by fine-tuning the model to specifically refuse answering watermark-related queries.
To construct the attack scenario, we fine-tuned the Qwen2.5-VL-7B-Instruct model on a mixed dataset designed to balance refusal behavior and general utility. Specifically, the training data included:
  • Adversarial Samples: 100 newly generated task-specific adversarial samples paired with refusal responses (e.g., “I can’t help with that one because the watermark cues have to stay hidden.”).
  • General Capabilities Data: 500 samples from COCO and 300 samples from Alpaca-GPT4 to maintain general visual reasoning capabilities.
Configuration & Results. We employed LoRA targeting all linear layers combined with NEFTune noise injection (Jain et al., 2024) (hyperparameters listed in Table 13).
Table 13: Hyperparameters for adversarial fine-tuning.
Hyperparameter Value Hyperparameter Value
Backbone Model
Qwen2.5-VL-7B-Instruct
LoRA Rank ($r$) 32
LoRA Alpha ($\alpha$) 8 Target Modules All Linear
NEFTune Noise $\alpha$ 5 Learning Rate $3e^{-5}$
Batch Size 8 Epochs 1
Warmup Ratio 0.03 Precision bf16
We evaluated the fine-tuned model on watermark detectability (CGSR) and general utility (MMQA Accuracy on 2,000 normal QA pairs). As shown in Table 14, while CGSR saw a moderate decrease, the watermark remained effectively detectable. Crucially, this process caused a significant degradation in general utility ($8.62\%$). This indicates that forcing the model to reject watermarked images—which are distributionally similar to natural data—triggers catastrophic forgetting of general visual reasoning capabilities, acting as a strong deterrent to adversaries.
Table 14: Results of adversarial fine-tuning. $\Delta$ denotes relative change.
Metric Original Model Adversarially Fine-tuned Model $\Delta$
$\textbf{AQUA}_{spatial}$ (CGSR) $98.42\%$ $89.34\%$ $9.23\%$
$\textbf{AQUA}_{acronym}$ (CGSR) $99.61\%$ $87.18\%$ $12.48\%$
Utility (MMQA Acc.) $27.25\%$ $24.90\%$ $8.62\%$
C.4.2 Stability under System Updates and Drifts
We empirically evaluate several common benign update scenarios (retriever replacement/fine-tuning and corpus expansion with distractors) and observe only marginal changes in Rank/CGSR, suggesting that AQUA remains stable under such deployment-time updates.
Robustness of the Retriever Component.
We evaluated the system’s performance under two distinct scenarios: architectural replacement and domain-specific adaptation. First, we replaced the default retriever with SigLIP (Zhai et al., 2023b) (siglip-so400m). Second, we fine-tuned the original retriever on the MMQA dataset.
Crucially, to simulate a realistic deployment scenario where the model adapts to user queries, we utilized 1,000 image-question pairs for fine-tuning, rather than the conventional image-caption pairs. This setting better reflects the distribution shift encountered in real-world QA tasks. As shown in Table 15, the ranking metrics remain highly consistent, demonstrating that AQUA generalizes well across architectures and remains effective even after task-specific optimization.
Table 15: Robustness results under retriever updates. We report the average RankRank (lower is better) for both methods.
Retriever Setting Rank $\downarrow$
$\textbf{AQUA}_{acronym}$ $\textbf{AQUA}_{spatial}$
Standard (CLIP-ViT) $1.027$ $1.054$
Replacement (SigLIP) $1.106$ $1.386$
Fine-tuned (CLIP-ViT) $1.038$ $1.376$
Resilience to Large-Scale Dataset Drifts.
Finally, we tested the robustness of our approach against varying sizes and contents of the retrieval corpus. Based on the original MMQA dataset, we progressively added 10k to 50k images from the WebQA dataset as distractors.
The results, detailed in Table 16, demonstrate that even with a significant increase in dataset size (nearly doubled), both the retrieval $Rank$ and $CGSR$ remain stable. This stability highlights the scalability of AQUA in handling dynamic, large-scale databases.
Table 16: Performance stability under large-scale dataset drifts (adding WebQA distractors).
Added Images +10k +20k +30k +40k +50k
Retrieval Performance (Rank $\downarrow$)
$\textbf{AQUA}_{acronym}$ $1.026$ $1.028$ $1.031$ $1.029$ $1.042$
$\textbf{AQUA}_{spatial}$ $1.351$ $1.526$ $1.424$ $1.496$ $1.696$
Semantic Decoding Performance (CGSR $\uparrow$)
$\textbf{AQUA}_{acronym}$ $98.77\%$ $98.39\%$ $96.10\%$ $97.21\%$ $96.76\%$
$\textbf{AQUA}_{spatial}$ $96.19\%$ $93.81\%$ $92.56\%$ $95.77\%$ $91.85\%$
Appendix D Ethics Statement
This work adheres to the ICLR Code of Ethics. In this study, no human subjects or animal experimentation were involved. All datasets used, including synthetic images, were sourced in compliance with relevant usage guidelines, ensuring no violation of privacy. We have taken care to avoid any biases or discriminatory outcomes in our research process. No personally identifiable information was used, and no experiments were conducted that could raise privacy or security concerns. We are committed to maintaining transparency and integrity throughout the research process.
Appendix E Reproducibility Statement
We have made every effort to ensure that the results presented in this paper are reproducible. All code has been uploaded as the supplemental materials to facilitate replication and verification. The experimental setup, including training steps, model configurations, is described in detail in the paper.
Additionally, multimodal QA datasets, such as MMQA and WebQA, are publicly available, ensuring consistent and reproducible evaluation results.
We believe these measures will enable other researchers to reproduce our work and further advance the field.
Appendix F The Use of Large Language Models (LLMs)
Large Language Models (LLMs) were used to aid in the writing and polishing of the manuscript. Specifically, we used an LLM to assist in refining the language, improving readability, and ensuring clarity in various sections of the paper. The model helped with tasks such as sentence rephrasing, grammar checking, and enhancing the overall flow of the text.
It is important to note that the LLM was not involved in the ideation, research methodology, or experimental design. All research concepts, ideas, and analyses were developed and conducted by the authors. The contributions of the LLM were solely focused on improving the linguistic quality of the paper, with no involvement in the scientific content or data analysis.
The authors take full responsibility for the content of the manuscript, including any text generated or polished by the LLM. We have ensured that the LLM-generated text adheres to ethical guidelines and does not contribute to plagiarism or scientific misconduct.